OpenStudy (kayders1997):

I think I did this right but I want to make sure solve the limit (1-sin0)/csc0 when 0 is approaching pi/2

OpenStudy (isaidavila):

What is your answer

OpenStudy (kayders1997):

0/1 or infinity @isaidavila

OpenStudy (mathmale):

0 does not approach pi/2. You'll need to use a variable, such as x or y, not the numeral 0. Please restate your problem (type it in again).

OpenStudy (kayders1997):

Hmmmm

OpenStudy (zarkon):

prob means \(\theta\)

OpenStudy (kayders1997):

I didn't how how to type that in

OpenStudy (kayders1997):

@mathmale

OpenStudy (mathmale):

You can use any variable you want: x, p, q, w, a, f, anything other than a constant such as 0.

OpenStudy (mathmale):

A variable can certainly approach pi/2; a constant such as 0 can never do that.

OpenStudy (kayders1997):

I did the symbol that Zarkon showed for the problem

OpenStudy (kayders1997):

@mathmale

OpenStudy (kayders1997):

@zepdrix

zepdrix (zepdrix):

whut :U

zepdrix (zepdrix):

\[\large\sf \lim_{\theta\to\pi/2}\frac{1-\sin \theta}{\csc \theta}\]So we're approaching \(\large\sf \dfrac{1-1}{1}\), ya?

OpenStudy (kayders1997):

Yes

zepdrix (zepdrix):

So we're approaching 0 divided by a number. What is \(\large\sf 0\div 4\) ?

OpenStudy (kayders1997):

0/4?

zepdrix (zepdrix):

Yes, what is 0 ..... divided by 4?

zepdrix (zepdrix):

If you take nothing, and divide it between 4 people how much does each person get?

OpenStudy (kayders1997):

0 I just don't understand where you got the 4

zepdrix (zepdrix):

It was an example :) So then \(\large\rm 0\div1=?\)

OpenStudy (kayders1997):

0 :) yay I got it

zepdrix (zepdrix):

yay

OpenStudy (kayders1997):

I feel proud lol

zepdrix (zepdrix):

yes go get a cookie :)

OpenStudy (kayders1997):

lol I had a cookie like 1 minute ago :D

zepdrix (zepdrix):

XD ok then stop there, dont let the diabedus get ya!

OpenStudy (kayders1997):

I love them tho haha

OpenStudy (kayders1997):

I hate stuff dealing with trig

zepdrix (zepdrix):

WHAT trig is bestest thing that ever happened to earth

OpenStudy (kayders1997):

Nope not in my mind I hated proving identities

OpenStudy (kayders1997):

I never closed this question?

zepdrix (zepdrix):

ya its closed

OpenStudy (kayders1997):

I never did lol